OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5 6 −6 All real numbers No solution

OpenStudy (anonymous):

no solution

OpenStudy (precal):

-3x-9=-3x-3-5 -3x-9=-3x-8 add 3x to both sides -9=-8 false no solution